Questor, Help Conduct the hypothesis test and provide the test statisisic, critical value and P-value, and state the conchusion. A person randomly selected 100 checks and recorded the cents portions of those checks T the indicated values. Use a 0.10 significance level he table below lists those cents portions categorized according to to test the claim that the four categories are equally likely. The person expected that many checks for r amounts would result in a disproportionately high frequency for the first category, but do the results support that expectation? Cents portion of check |0-24 Number 25-49 23 14 75.99 50-74 30 The test statisticis(Type an integer of a decimal) The critical value is(Round to three decimal places as needed) The P-value is □ (Round to four decimal places as needed ) State the conclusion Ho. There suficient evidence to warrant rojection of the caim that the four catlegates are equalykely The ruo to support the expectation that the frequency for the first category ia dioproportionately high
Answer #1

applying chi square goodness of fit test:

           relative observed Expected residual Chi square
category frequency Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
0-24 0.250 30.000 25.00 1.00 1.000
25-49 0.250 23.000 25.00 -0.40 0.160
50-74 0.250 26.000 25.00 0.20 0.040
75-99 0.250 21.000 25.00 -0.80 0.640
total 1.000 100 100 1.840

test statistic =1.840

critical value=6.251

p value =0.6063

fail to reject Ho ; there is not sufficient .........the result does not support the expectation,,,,,,,,,,